# Particle Trajectory Representation Learning with Masked Point Modeling

### v1

@@ 2 Related Work
Masked modeling: At the heart of any self-supervised learning framework is the pretext task, i.e., the task under which the model will learn good representations of the data. For our purposes, we focus on two types of pretext tasks: masked modeling (Devlin, 2018) and contrastive learning (Chen et al., 2020b; He et al., 2020; Oord et al., 2018; Tian et al., 2020a, b). The masked modeling task involved masking a part of the training sample, and tasking the model with predicting what it did not see. The success of masked image (He et al., 2021) and language (Devlin, 2018) modeling prompted a number of works to propose masked modeling frameworks for point cloud understanding. Point-MAE (Pang et al., 2022) and Point-BERT (Yu et al., 2022) are perhaps the most popular frameworks, of which many models today use as a baseline for comparison. Point-MAE extended image masked autoencoders (MAEs) (He et al., 2021) to the point-cloud regime by explicitly reconstructing point cloud groupings using the Chamfer distance. Point-BERT extended BERT (Devlin, 2018) pre-training to point clouds. In this work, we start with the Point-MAE architecture, as it is the most conceptually simple and does not require pre-training a point cloud tokenizer as in Point-BERT. In contrast, the contrastive learning task involves maximizing the similarity between augmentations of the same training sample while minimizing the similarity of other samples. Though extremely successful in image tasks, we opt to explore self-distillation for particle trajectory representation learning in future works, just focusing on whether SSL for LArTPC images is possible in the first place.
@@ 3 Dataset
The training set consists of 1,033,377 events containing at least 1024 points. A smaller dataset of approximately 30 events (15,000 tokens) from a cache of 10,471 held-out events are used for validation metrics for pre-training (see below), and another 30 events from the same cache are used for testing. For fine-tuning, we use a variable number of events from the training set, and use all 10,471 held-out events in the validation set for evaluation.
@@ 4.1.1 Patch grouping
The resulting groups form a variable number $G$ of point groups, each containing a variable number $K$ of points. This approach approximates the effect of voxelizing the volume spanned by all points with a voxel pitch equal to the sphere diameter, while inherently reducing aliasing artifacts. Unlike FPS+$k$-NN, which lacks control over patch overlap, C-NMS ensures a tunable overlap between patches. We argue that $k$-NN grouping post-C-NMS is worse than a ball query, as the latter naturally constrains the domain of points within a canonical radius, effectively mitigating outlier inclusion within patches. Additional commentary on this approach can be found in Appendix B.
@@
In traditional FPS+{$k$-NN, ball query}, researchers often employ a fixed number of patches and a fixed number of points per patch (e.g., someone pre-training on ShapeNet (Chang et al., 2015) may produce 128 patches with 32 points each). By contrast, C-NMS dynamically determines both the number of tokens $G$ and the number of points per token $K$ based on two parameters: the sphere radius $r$ and the overlap factor $f$. Algorithm 1 provides a pseudocode implementation of C-NMS. The computational overhead of the additional NMS step involves an extra ball query on seed points, which can be efficiently executed using CUDA-accelerated libraries such as pytorch3d (Ravi et al., 2020), but is certainly non-negligible.
@@ 5.1 Pre-training
Table 1: SVM Validation Results. We present final SVM validation F1 scores for each class—where F1 is the harmonic mean of precision and recall—and an overall mean across classes. Bold entries highlight the best results.
@@
PointMAE 100k 0.852 0.990 0.970 0.436 0.325
PointMAE 500k 0.886 0.993 0.967 0.572 0.342
PoLAr-MAE 500k 0.895 0.994 0.977 0.518 0.440
@@ Appendix C Hyperparameters
We provide detailed hyperparameter information in the following sections. The full implementation of these models, as well as training code, can be found at <url>.
@@ Appendix F Failures
  • Auxiliary Energy Reconstruction Task. We introduced an auxiliary task to reconstruct energy values for masked-out groups of points. Masked 3D points were embedded, concatenated with the regressed mask tokens from the decoder, and passed through an additional regression head to predict energies for each point. To achieve this, we modified the mini-PointNet to be permutation equivariant rather than invariant by adding learned point ordering embeddings. These embeddings were computed using positional encoding (Vaswani, 2017) combined with a one-layer MLP and added to the uplifted point features after the first shared MLP layer in the mini-PointNet. Although the model successfully reconstructed the energies, this task did not lead to any significant improvement in semantic type discrimination.

### v2

@@ 2 Related Work
Masked modeling: At the heart of any self-supervised learning framework is the pretext task, i.e., the task under which the model will learn good representations of the data. For our purposes, we focus on two types of pretext tasks: masked modeling (Devlin, 2018) and contrastive learning (Chen et al., 2020b; He et al., 2020; Oord et al., 2018; Tian et al., 2020a, b). The masked modeling task involved masking a part of the training sample, and tasking the model with predicting what it did not see. The success of masked image (He et al., 2021) and language (Devlin, 2018) modeling prompted a number of works to propose masked modeling frameworks for point cloud understanding. Point-MAE (Pang et al., 2022) and Point-BERT (Yu et al., 2022) are perhaps the most popular frameworks, of which many models today use as a baseline for comparison. Point-MAE extended image masked autoencoders (MAEs) (He et al., 2021) to the point-cloud regime by explicitly reconstructing point cloud groupings using the Chamfer distance. Point-BERT extended BERT (Devlin, 2018) pre-training to point clouds. In this work, we start with the Point-MAE architecture, as it is the most conceptually simple and does not require pre-training a point cloud tokenizer as in Point-BERT. In contrast, the contrastive learning task involves maximizing the similarity between augmentations of the same training sample while minimizing the similarity of other samples. Though extremely successful in image tasks, we opt to explore contrastive learning approaches for particle trajectory representation learning in future works, just focusing on whether SSL for LArTPC images is possible in the first place.
@@ 3 Dataset
The training set consists of 1,033,377 events containing at least 1024 points. A smaller dataset of approximately 30 events (15,000 tokens) from a cache of 10,471 held-out events are used for validation metrics for pre-training, and another 30 events from the same cache are used for testing. For fine-tuning, we use a variable number of events from the training set, and use all 10,471 held-out events in the validation set for evaluation.
@@ 4.1.1 Patch grouping
The resulting groups form a variable number $G$ of point groups, each containing a variable number $K$ of points. This approach approximates the effect of voxelizing the volume spanned by all points with a voxel pitch equal to the sphere diameter, while inherently reducing aliasing artifacts. Unlike FPS+$k$-NN, which lacks control over patch overlap, C-NMS ensures a tunable overlap between patches. We argue that $k$-NN grouping post-C-NMS is worse than a ball query, as the latter naturally constrains the domain of points within a canonical radius, effectively mitigating outlier inclusion within patches. Additional commentary on this approach, including the pseudocode implementation, can be found in Appendix B.
@@
In traditional FPS+{$k$-NN, ball query}, researchers often employ a fixed number of patches and a fixed number of points per patch (e.g., someone pre-training on ShapeNet (Chang et al., 2015) may produce 128 patches with 32 points each). By contrast, C-NMS dynamically determines both the number of tokens $G$ and the number of points per token $K$ based on two parameters: the sphere radius $r$ and the overlap factor $f$. The computational overhead of the additional NMS step involves an extra ball query on seed points, which can be efficiently executed using CUDA-accelerated libraries such as pytorch3d (Ravi et al., 2020), but is certainly non-negligible.
@@ 5.1 Pre-training
Table 1: SVM Validation Results. We present final SVM validation F1 scores for each class—where F1 is the harmonic mean of precision and recall—and an overall mean across classes. The best results are highlighted.
@@
PointMAE 100k 0.680 0.990 0.970 0.436 0.325
PointMAE 500k 0.719 0.993 0.967 0.572 0.342
PoLAr-MAE 500k 0.732 0.994 0.977 0.518 0.440
@@ Appendix C Hyperparameters
We provide detailed hyperparameter information in the following sections. The full implementation of these models, as well as training code, can be found at https://github.com/DeepLearnPhysics/PoLAr-MAE/.